# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ Abstract
Conformal prediction builds marginally valid prediction intervals which cover the unknown outcome of a randomly drawn new test point with a prescribed probability. In practice, a common scenario is that, after seeing the test unit(s), practitioners decide which test unit(s) to focus on in a data-driven manner, and wish to quantify the uncertainty for the focal unit(s). In such cases, marginally valid prediction intervals for these focal units can be misleading due to selection bias. This paper presents a general framework for constructing a prediction set with finite-sample exact coverage conditional on the unit being selected. Its general form works for arbitrary selection rules, and generalizes Mondrian Conformal Prediction to multiple test units and non-equivariant classifiers. We then work out computationally efficient implementation of our framework for a number of realistic selection rules, including top-K selection, optimization-based selection, selection based on conformal p-values, and selection based on properties of preliminary conformal prediction sets. The performance of our methods is demonstrated via applications in drug discovery and health risk prediction.
@@ 1 Introduction
Conformal prediction is a versatile framework for quantifying the uncertainty of any black-box prediction model, by issuing a prediction set that covers an unknown outcome with a prescribed probability. Formally, suppose the task is to predict an outcome $Y\in{\mathcal{Y}}$ based on features $X\in{\mathcal{X}}$. Given a set of calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the features of a new test point $X_{n+1}$, conformal prediction builds upon a given prediction model and delivers a prediction set $\widehat{C}_{\alpha,n+1}\subseteq{\mathcal{Y}}$ at level $\alpha\in(0,1)$, which obeys
@@
With this finite-sample, distribution-free guarantee, the conformal prediction set $\widehat{C}_{\alpha,n+1}$ describes a range of plausible values the unknown outcome $Y_{n+1}$ can take, thereby expected to inform downstream decision-making based on any black-box prediction model. With such a promise, methods for issuing marginally-valid – in the sense of (1) – prediction sets have been developed for various problems; see e.g., Angelopoulos and Bates, (2021) for a review.
@@
In many downstream applications, however, people are often only interested in a selective subset of units. For example, practitioners may act upon a unit if it exhibits an interesting property (Olsson et al.,, 2022; Sokol et al.,, 2024; Levitskaya,, 2023); in some cases, practitioners may want to focus on a subset suggested by some complicated deta-dependent decision process such as resource optimization (Svensson et al.,, 2018; Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). It would be misleading if the prediction sets fail to deliver the promised coverage guarantee for the selected units. Let us discuss a few applications where such cases may arise.
@@
  • In drug discovery, an important task is to predict the binding affinity of a drug candidate to a target to inform drug prioritization (Laghuvarapu et al.,, 2024). Scientists may only focus on drug candidates with highest predicted affinites, or selected by an false discovery rate (FDR)-controlling procedure (Jin and Candès, 2023b, ), or whose prediction sets only cover high values (Svensson et al.,, 2017), or that optimize resource usage (Svensson et al.,, 2018). It will be misleading if more than an $\alpha$-fraction of prediction sets for the selected drugs fail to cover the actual binding affinities.
@@
  • In business decision-making, companies may take different inventory decisions based on whether a conformal prediction set suggests a strong demand or a weak demand (Levitskaya,, 2023). Similarly, it will be problematic if the average coverage among strong-demand predictions is below $1-\alpha$.
@@
  • In disease diagnosisOlsson et al., (2022) suggest human intervention if the prediction set for a disease status is too large (this implicitly declares confidence in small prediction sets; see similar ideas in Ren et al., (2023); Sokol et al., (2024)). However, it would be concerning if more than an $\alpha$-fraction of small-sized prediction sets (“approved” as confident) miss the true disease status.
@@
  • In healthcare management, clinical decision makers may send patients to different healthcare categories after optimizing a performance measure (such as waiting time) subject to certain constraints, such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). It is thus natural to only include a data-dependent subset of patients in downstream uncertainty quantification.
@@
In all these cases, it is highly desirable that a prediction set should cover the unknown outcome for a unit of interest with a prescribed probability. This motivates a stronger, selection-conditional guarantee. Supposing there are $m\geq 1$ test units $\{X_{n+j}\}_{j=1}^{m}$, we shall aim for
@@
where ${\mathcal{D}}_{\textnormal{calib}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is the calibration data, ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ is the test data, and ${\mathcal{S}}(\cdot,\cdot)\subseteq[m]:=\{1,\dots,m\}$ is a data-driven process to decide which units are of interest. In particular, we focus on the case where $\widehat{C}_{\alpha,n+j}$ is constructed after the selection set ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is determined. This is similar to post-selection inference (Lee et al.,, 2016; Markovic et al.,, 2017; Tibshirani et al.,, 2018), yet we study a predictive inference problem; see Section 1.2 for more discussion.
@@
A prediction set with marginal validity (1) does not necessarily cover an unknown outcome conditional on being of interest as (2). Such selection issue has caught recent attention in the literature of predictive inference. Through analysis of a real drug discovery dataset, Jin and Candès, 2023b (, Section 1) demonstrate that more than $30\%$ of seemingly promising prediction sets (in the sense that $\widehat{C}_{\alpha}(X_{n+1})=\{1\}$) miss the actual outcomes for $\alpha=0.01$, i.e., with a marginal coverage of $99\%$. We shall see more examples of this issue in our numerical experiments under various selection processes. Focused on downstream decisions, Jin and Candès, 2023b ; Jin and Candès, 2023a develop methods that select units with large outcomes while guranteeing control of FDR. For uncertainty quantification, Bao et al., (2024) propose methods to construct prediction sets for selected test units with (approximately) false coverage rate (FCR) control. Their methods only achieve exact guarantees for a limited class of selection rules such as top-$K$ selection, while we are to present a general framework that works for arbitrary selection rules. In addition, FCR control and selection-conditional coverage (2) are related but distinct error notions, which we discuss in Section 2.2.
@@
We also note that the selection-conditional coverage (2) may be implied by other stronger notions such as conditional coverage: $\mathbb{P}(Y_{n+1}\in\widehat{C}_{\alpha,n+1}{\,|\,}X_{n+1}=x)\geq 1-\alpha$ for $\mathbb{P}$-almost all $x\in{\mathcal{X}}$. Conditional coverage implies (2) if $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are mutually independent and the selection rule only depends on the features. However, it is not achievable by finite-length prediction intervals in the absence of distributional assumptions (Foygel Barber et al.,, 2021), and the underlying selection rule may be more complicated. In contrast, we may view (2) as a “relaxed” version of conditional coverage that is both achievable and relevant for practical use.
1.1 Overview of contributions
In this paper, we propose a unified framework, named JOint Mondrian Conformal Inference (JOMI), that builds valid prediction sets for focal units picked by any data-dependent selection process. To this end, we make the following contributions.
@@ 1.1 Overview of contributions
In Section 3, we introduce the general formulation of JOMI, whose crux is to find a data-dependent “reference” set of calibration data that are exchangeable conditional on the selection event. In this sense, our method can be viewed as a generalization of Mondrian Conformal Prediction (Vovk et al.,, 2005). We prove that for exchangeable data, the prediction set $\widehat{C}_{\alpha,n+j}$ produced by JOMI covers $Y_{n+j}$ with probability at least $1-\alpha$ conditional on a selection event. The selection event can be “test unit $j$ is selected”, which leads to (2). It can also be more granular, such as “test unit $j$ is selected, and there are $k$ selected test units”. Our framework works for any selection rule ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that is permutation-invariant to ${\mathcal{D}}_{\textnormal{calib}}$, without any modeling assumptions on the data generating process.
In Section 4, we work out efficient implementation of JOMI for a number of selection processes that may be of practical interest, including:
@@
  • Covariate-dependent selection. When the selection rule do not involve $\{Y_{i}\}_{i=1}^{n}$, the computation complexity of our method is (at most) $O(mn)$ times that of the selection rule. This imples efficient implementation for a wide range of problems, including various forms of top-K selection and selecting based on complicated constrained optimization programs. As a special case, we recover the method of Bao et al., (2024) for top-K selection among test data, and yield valid solutions to other ranking-based selection methods attempted in their work.
@@
  • Conformal p-value-based selection. We also derive efficient implementation for selection rules based on conformal p-values with both fixed and data-dependent thresholds. For instance, our prediction set is finite-sample exactly valid for units selected by Conformalized Selection (Jin and Candès, 2023b, ), which is studied in Bao et al., (2024) with approximate guarantee.
@@
  • Selection based on preliminary conformal prediction set. In addition, we derive a general, efficient instantiation when unit(s) whose marginal prediction sets demonstrate an interesting property are selected, such as being of a short length, or having a lower bound above a threshold. Our methods can be useful for re-calibrating the uncertainty quantification of such seemingly promising units.
@@
Finally, we demonstrate the applications of our methods via a few realistic selection rules that may occur in drug discovery (Section 5) and health risk prediction (Section 6). Our results show that marginally valid prediction sets may undercover or overcover, while our methods lead to valid coverage.
1.2 Related work
Our methods extend conformal prediction, which is a general framework for building marginally valid prediction intervals (Vovk et al.,, 2005; Lei et al.,, 2018; Romano et al.,, 2019, 2020). Compared with vanilla conformal prediction, we aim at a new goal, which is to achieve selection-conditional coverage for a unit of interest, instead of marginally for an “average” unit. Particularly related to our framework is Mondrian Conformal Prediction (Vovk et al.,, 2003, MCP), which uses a subset of calibration data to build a prediction interval for coverage conditional on a pre-determined or permutation-invariant class membership. However, achieving selection-conditional coverage is a new goal with distinct motivations, and our techniques and scopes of application are significantly different (see Remark 1).
The selection-conditional guarantee we seek for is closely related to post-selection inference (POSI) literature (Berk et al.,, 2013; Tibshirani et al.,, 2016; Chernozhukov et al.,, 2015) and in particular conditional selective inference (Lee et al.,, 2016; Markovic et al.,, 2017; Reid et al.,, 2017; Tibshirani et al.,, 2018; Kivaranovic and Leeb,, 2020; McCloskey,, 2024). In POSI, one important task is to build confidence intervals for a selected subset of model parameters conditional on the selection event, which is close to our goal of selection-conditional coverage guarantee. Methods for this goal usually leverage specific problem structure such as linearity and distribution of the estimators to derive the confidence intervals (Lee et al.,, 2016; Tian and Taylor,, 2017; Andrews et al.,, 2024; Liu,, 2023). As we focus on predictive inference (we have units instead of parameters), we leverage the exchangeability across units to achieve valid selection-conditional coverage, hence our methods are very different.
Another related error notion in the selective inference literature is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005); we compare them in more details in Section 2.2. Methods for achieving FCR control have been proposed for various settings (Benjamini and Yekutieli,, 2005; Weinstein et al.,, 2013; Weinstein and Ramdas,, 2020; Zhao and Cui,, 2020). For conformal prediction, Bao et al., (2024) propose prediction sets for selected units while achieving (approximate) FCR control. However, their methods only work for multiple test units. In addition, they achieve exact guarantee only for selection rules that are invariant to permutations of calibration and test units and top-$K$ selection (where $K$ test units with the highest predictions are selected); we will show that when applying JOMI to these problems, we yield the same prediction sets as theirs. Meanwhile, our framework leads to exact guarantee for some other problems studied in Bao et al., (2024) with approximate FCR control. Our framework also tackles many more settings with complicated selection rules that cannot be handled by theirs.
Selective inference issues in conformal prediction or predictive inference have caught recent attention. Most of existing works study multiple testing, i.e., selecting individuals that obey some properties with false discovery rate (FDR) control, including selecting outliers (Bates et al.,, 2023; Marandon et al.,, 2022; Liang et al.,, 2022) and selecting high-quality data (Jin and Candès, 2023b, ; Jin and Candès, 2023a, ). The construction of conformal prediction sets is an orthogonal direction, and our methods also work for selection based on such procedures (we develop an efficient method in Section 4.2.2).
Under a similar name, Sarkar and Kuchibhotla, (2023) proposed “post-selection inference for conformal prediction”, but they mainly focus on selecting the confidence level $\alpha\in(0,1)$ instead of selecting the units. In addition, they fall within the “simultaneous inference” strand in POSI so that their inference is valid simultaneously for all confidence levels, other than conditional on a selection event. For these reasons, our methods and guarantees are quite different.

### v3

@@ Abstract
Conformal prediction builds marginally valid prediction intervals that cover the unknown outcome of a randomly drawn test point with a prescribed probability. However, in practice, data-driven methods are often used to identify specific test unit(s) of interest, requiring uncertainty quantification tailored to these focal units. In such cases, marginally valid conformal prediction intervals may fail to provide valid coverage for the focal unit(s) due to selection bias. This paper presents a general framework for constructing a prediction set with finite-sample exact coverage, conditional on the unit being selected by a given procedure. The general form of our method accommodates arbitrary selection rules that are invariant to the permutation of the calibration units, and generalizes Mondrian Conformal Prediction to multiple test units and non-equivariant classifiers. We also work out computationally efficient implementation of our framework for a number of realistic selection rules, including top-K selection, optimization-based selection, selection based on conformal p-values, and selection based on properties of preliminary conformal prediction sets. The performance of our methods is demonstrated via applications in drug discovery and health risk prediction.
@@ 1 Introduction
Conformal prediction is a versatile framework for quantifying the uncertainty of any black-box prediction model, by issuing a prediction set that covers the unknown outcome with a prescribed probability. Formally, suppose the task is to predict an outcome $Y\in{\mathcal{Y}}$ based on features $X\in{\mathcal{X}}$. Given a set of calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the features of a new test point $X_{n+1}$, conformal prediction builds upon a given prediction model and delivers a prediction set $\widehat{C}_{\alpha,n+1}\subseteq{\mathcal{Y}}$ at level $\alpha\in(0,1)$, which obeys
@@
With this finite-sample, distribution-free guarantee, the conformal prediction set $\widehat{C}_{\alpha,n+1}$ describes a range of plausible values the unknown outcome $Y_{n+1}$ may take, thereby expected to inform downstream decision-making based on the black-box prediction model. With such a promise, methods for constructing marginally valid – in the sense of (1) – prediction sets have been developed for various problems; see e.g., Angelopoulos and Bates, (2021) for a review.
@@
In many downstream applications, however, people are often only interested in a selective subset of units. For example, practitioners may only act upon a unit if it exhibits an interesting property (Olsson et al.,, 2022; Sokol et al.,, 2024; Levitskaya,, 2023), or focus only on a subset of test units picked by a complicated data-dependent process such as resource optimization (Svensson et al.,, 2018; Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). It would be misleading for the practitioners if the prediction sets fail to deliver the promised coverage guarantee for the selected unit(s). Let us discuss a few applications where such cases may arise.
@@
  • In drug discovery, an important task is to predict the binding affinity of a drug candidate to a disease target, which informs subsequent drug prioritization (Laghuvarapu et al.,, 2024). Among many drug candidates, scientists may only focus on those with highest predicted affinities, or those selected by a false discovery rate (FDR)-controlling procedure (Jin and Candès, 2023b, ), or whose prediction sets only cover high values (Svensson et al.,, 2017), or those optimizing resource usage (Svensson et al.,, 2018). It may lead to a waste of resources if the prediction sets for the selected drugs fail to cover the actual binding affinities with an exceedingly high chance.
@@
  • In business decision-making, companies may take different inventory decisions based on whether a conformal prediction set suggests a strong demand or a weak demand (Levitskaya,, 2023). Similarly, it will be problematic if a strong-demand prediction cannot cover with at least $1-\alpha$ of the time.
@@
  • In disease diagnosisOlsson et al., (2022) suggest human intervention if the prediction set for a disease status is too large (this implicitly declares confidence in small prediction sets; see similar ideas in Ren et al., (2023); Sokol et al., (2024)). However, it would be concerning if, with more than a chance $\alpha$, the small-sized prediction sets – “approved” as confident – miss the true disease status.
@@
  • In healthcare management, patients may be sent to different healthcare categories based a program that optimizes some performance measure (such as waiting time) subject to certain constraints, such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014). The subset of patients in each category is therefore data-dependent.
@@
In all these examples, it is highly desirable that a prediction set should cover the unknown outcome for a unit of interest with a prescribed probability. This motivates a stronger, selection-conditional guarantee. Supposing there are $m\geq 1$ test units ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$, we aim for
@@
where ${\mathcal{D}}_{\textnormal{calib}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is the calibration data, and ${\mathcal{S}}(\cdot,\cdot)$ is a data-driven process to decide the units of interest, which maps the calibration and test data to a subset of $[m]:=\{1,\dots,m\}$. Our target is similar in spirit to post-selection inference (Lee et al.,, 2016; Markovic et al.,, 2017; Tibshirani et al.,, 2018), but we consider predictive inference settings and develop distinct techniques; see Supplementary Section A.1 for more discussion. Throughout, we focus on settings where the prediction sets are constructed after ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is determined. By separating the selection process from the (post hoc) uncertainty quantification step, we leave the freedom of defining selection rules to the practitioners.
@@
A prediction set with marginal validity (1) does not necessarily cover an unknown outcome conditional on being of interest as in (2). Such a selection issue has recently been raised in the literature of predictive inference: through analysis of a real drug discovery dataset, Jin and Candès, 2023b (, Section 1) demonstrate that more than $30\%$ of seemingly promising prediction sets (in the sense that $\widehat{C}_{\alpha,n+j}=\{\text{active}\}$) miss the actual outcomes when the nominal marginal miscoverage rate is $\alpha=0.01$. We shall see more examples of this issue in our numerical experiments with various selection processes. New techniques are therefore needed for constructing prediction sets achieving (2).
1.1 Exchangeability via reference sets
Accounting for selection in conformal prediction is a delicate task since it breaks exchangeability. The core of conformal prediction is to leverage the exchangeability among the calibration and test data, such that their “prediction residuals”, referred to as nonconformity scores, are comparable in distribution (see Section 3.1 for more details). The calibration scores therefore inform the magnitude of uncertainty in a new test point (Vovk et al.,, 2005). However, given a selection event, the calibration data are no longer exchangeable with the test point, leading to violation of the coverage guarantee mentioned above. In general, the selection-conditional distributions of these scores are complex since the selection event can be highly data-dependent.
Such a challenge motivates our new framework, named JOint Mondrian Conformal Inference (JOMI), which builds prediction sets that achieve selection-conditional coverage, with (2) as a special case. As visualized in Figure 1, our key idea is to find a “reference set” — a data-dependent subset of calibration data that remain exchangeable with respect to the new test point conditional on the selection event. This reference set thus provides calibrated quantification of uncertainty for a selected unit. The mechanism we devise accommodates arbitrary selection rules ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that are invariant to permutations of data in ${\mathcal{D}}_{\textnormal{calib}}$, and can be computed efficiently for a wide range of commonly used selection rules.
Figure 1: Visualization of the intuition behind the reference set. (a) Marginally, the calibration data (black) are exchangeable with respect to the test point (blue). (b) The calibration data are not exchangeable with respect to the test point (shaded blue) given a selection event. (c) We find calibration data which, when posited as a “test point”, would lead to the same selection event. (d) The reference set consists of calibration data that are exchangeable with respect to the test point given selection, and we use them to construct JOMI prediction sets.
@@ 1.1 Exchangeability via reference sets
We also note that the selection-conditional coverage (2) may be implied by other stronger notions such as conditional coverage: $\mathbb{P}(Y_{n+1}\in\widehat{C}_{\alpha,n+1}{\,|\,}X_{n+1}=x)\geq 1-\alpha$ for $\mathbb{P}$-almost all $x\in{\mathcal{X}}$. Conditional coverage implies (2) if $\{(X_{i},Y_{i})\}_{i=1}^{n+1}$ are mutually independent and the selection rule only depends on test features. However, it is not achievable by finite-length prediction intervals without distributional assumptions (Vovk et al.,, 2005; Barber et al.,, 2021), and practical selection rules may also depend on other information. Given these considerations, we may view (2) as a “relaxed” version of conditional coverage that is both achievable and relevant for practical use.
1.2 Preview of contributions
@@
In Section 3, we introduce the general formulation of JOMI, including the construction of reference set and its use in deriving a selection-conditionally valid prediction set. We prove that under exchangeability, the prediction set $\widehat{C}_{\alpha,n+j}$ produced by JOMI covers $Y_{n+j}$ with probability at least $1-\alpha$ conditional on a selection event. The selection event can be “test unit $j$ is selected”, which leads to (2); it can also be more granular, such as “test unit $j$ is selected, and there are $k$ selected test units”. This framework is valid for any selection rule ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ that is permutation-invariant to ${\mathcal{D}}_{\textnormal{calib}}$, without any modeling assumptions on the data generating process.
In Section 4, we study the computational aspect of JOMI. We show that when $|{\mathcal{Y}}|<\infty$, our method can be computed with a worst-case complexity of $O(|{\mathcal{Y}}|mn)$ times that of the selection rule. Moreover, for general continuous outcome space ${\mathcal{Y}}$, we work out efficient implementation of JOMI for a number of selection processes that may be of practical interest, including:
@@
  • Covariate-dependent selection. When the selection rule does not involve $\{Y_{i}\}_{i=1}^{n}$, the computation complexity of our generic method is (at most) $O(mn)$ times that of the selection rule. This implies efficient implementation for a wide range of problems, including various forms of top-K selection (for which the computation can be further reduced to $O(m+n)$) and selection based on complicated constrained optimization programs. As a special case, we recover the method of Bao et al., (2024) for top-K selection among test data, and provide valid solutions to other ranking-based selection methods attempted in their work.
@@
  • Conformal p-value-based selection. We derive efficient implementation for a class of selection rules based on thresholding conformal p-values with “stopping time-type” cutoffs. For instance, our prediction set is finite-sample exactly valid for units selected by the Conformal Selection method (Jin and Candès, 2023b, ), which is studied in Bao et al., (2024) with approximate FCR guarantee.
@@
  • Selection based on preliminary conformal prediction sets. In addition, we derive a general, efficient instantiation when selecting units whose marginal prediction sets demonstrate certain interesting properties, such as being of a short length, or having a lower bound above some threshold. Our method can be useful for re-calibrating the uncertainty quantification of such seemingly promising units.
@@
Finally, we demonstrate the application of our methods via several realistic selection rules that may occur in drug discovery (Section 5) and health risk prediction (Section 6). Our results show that marginal prediction sets may undercover or overcover the selected units, while our methods always achieve the promised coverage guarantees.
Due to space constraints, we delegate a comprehensive literature review to Supplementary Section A.1. We end this section with some useful notations.
Notations.
For a positive integer $n\in\mathbb{N}^{+}$, write $[n]:=\{1,\dots,n\}$. We write the data pair as $Z=(X,Y)$, so that the calibration data is ${\mathcal{D}}_{\textnormal{calib}}=\{Z_{i}\}_{i=1}^{n}$. For any $j\in[m]$, we define the augmented calibration set as ${\mathcal{D}}_{j}={\mathcal{D}}_{\textnormal{calib}}\cup\{Z_{n+j}\}$ and the remaining test set as ${\mathcal{D}}^{c}_{j}={\mathcal{D}}_{\textnormal{test}}\backslash\{X_{n+j}\}$. The unordered set of ${\mathcal{D}}_{j}$ is denoted $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$, which provides the order statistics.